# Self-Guidance: Boosting Flow and Diffusion Generation on Their Own

### v4

@@ 5 Experiments
Table 1: Quantitative results of different Stable Diffusion models and Flux with various guidance. FID, Clip Score and Aesthetic Score are evaluated on the MS-COCO 2017 validation subset. Bolded values highlight the best performance for each metric in each method.
@@
Flux.1[29] 67.0163 0.9580 34.1976 4.5448 0.3080 29.74 22.9802 1.0971
+SG 66.4360 0.9616 34.0561 4.6634 0.3084 28.59 22.9960 1.1046
Table 2: Quantitative results of SG Performances in Specific Generative Domains compared with CFG baseline Flux.1 [29].
5.1 Implementation Details
@@
We apply our Self-Guidance on current state-of-the-art text-to-image and text-to-video models. For text-to-image, we consider both a classic diffusion model, Stable Diffusion 1.4 [52], and current state-of-the-art flow-based models, Stable-Diffusion 3 series [8, 2] and Flux.1 [29]. For text-to-video, we consider CogVideoX-5B [67], an open-source video generation model trained with flow matching.
To facilitate a fair comparison, we consistently evaluate the model with the same number of steps in inference: DDIM Solver with 50 steps for SD 1.4, Euler with 28 steps for SD 3 series, and FLUX. The guidance scales for CFG and PAG are kept the same as the original settings in the corresponding papers. We will elaborate in Appendix C.3.
@@
5.2 Quantitative Comparisons
5.2.1 Text-to-Image Generation
Table 1 shows the effect of using SG itself on the three SD models. First of all, by incorporating SG itself on SD 1.4, we achieve an impressive 8-point (50.7602$\to$42.7362) improvement in the FID. Similar improvements appear in other metrics and models as well. This indicates that SG itself is able to improve the quality of generated images.
SG is specifically designed to address the artifact issues in CFG and PAG. As shown in Table 1, incorporating SG with CFG leads to consistent improvements in HPS across all models, demonstrating that SG better aligns with human preferences. Moreover, our approach remains competitive with PAG in both CLIP and Aesthetic Score.
Then, we find that combining SG with CFG and PAG will have the best performance over using them alone. In SD 1.4 and SD 3, this integration yields significant qualitative improvements, as detailed in the Appendix A.1. The results show that adding SG not only enhances fine-grained details, prompt alignment, and error correction but also achieves new SoTA better performance across multiple metrics (24.65 HPS v2.1 for SD 1.4, 25.28 FID for SD 3, 0.3132 CLIP Score in SD 3.5, and 5.7845 Aesthetic Score in Flux).
In addition, it is surprising to find that SG-prev outperformed both compared baselines and the SG model, particularly in terms of FID and CLIP Score. This is achieved by applying SG-prev after the diffusion time (not step) $t$ decreases below 500, because the earlier steps generate too noisy images with clueless SG signals. This demonstrates that the SG-prev successfully boosts the diffusion models without incurring any additional cost.
5.2.2 Text-to-Video Generation
Table 3 provides quantitative evaluation results across four key dimensions in VBench [19]. The results indicate that adding SG effectively improves the metrics in human action (0.8320$\to$0.8450), and multiple objects (0.4619$\to$0.4703). SG-prev also outperforms the baseline across all four metrics under the same sampling time, but slightly lags behind SG that doubles sampling time. After having a close look at the generated examples in Figure 10, we find that SG also effectively eliminates artifact problems (broken arms and legs or misplaced limbs) in video generation tasks and enhances text-video consistency. Due to this comparison, we may conclude that SG is also effective in improving video generation quality. We will provide more video examples in Appendix A.2.
@@ 5.2.2 Text-to-Video Generation
CogVideoX-5B[67] 0.8320 0.2526 0.4619 0.2331
+ SG 0.8450 0.2553 0.4703 0.2339
+SG-prev 0.8376 0.2530 0.4774 0.2333
Table 3: Quantitative results of CFG based text-to-video generation model CogVideoX with our SG and SG-prev. All scores are evaluated on the Standard Prompt Suite of VBench[19]. Bolded values highlight the best performance for each metric.
Figure 6: Ablation study of the Guidance Scale and Shift Scale.
5.3 Guidance Performances in Specific Domains
Through our experiments and analysis (e.g. Figure 2), we observed that SG is particularly effective when handling some long-standing problems in image generation. More specifically, SG can effectively generate hands with an exact correct number of fingers with natural shapes (Part 1), and remove artifacts like redundant arms or hands (Part 2). It also enhances Flux’s comprehension of textual input (Part 4), effectively eliminating extraneous objects. This allows the main subject of the image to stand out while the background is appropriately blurred(Part 3).
@@ 5.3 Guidance Performances in Specific Domains
As shown in Table 2, SG achieves significant results in fine-grained generation tasks, such as hand and face generation, with improvements in FID-H (67.0163 to 66.4360) and FID-F (34.1976 to 34.0561). Additionally, the increase in PickScore and ImageReward indicates that our SG is more adept at producing higher-quality samples, thus minimizing generation errors, as illustrated in Part 4 of Figure 2.
@@
5.4 Ablation Studies
5.4.1 Guidance Scale
@@ 5.4.2 Shift Scale
The shift scale $\delta(t)$ represents the extent to which the noise level deviates from the current level. A simple way is to set it to a constant independent of $t$. The larger the shift scale is, the more it could contrast with the current noise level to suppress the artifacts. However, too large shift may damage the desired patterns that should be preserved in the current sampling density, and thus a suitable shift scale should be adopted. Figure 6 shows how the FID and HPS score change as the shift scale increases from 0 to 50 (within 1,000 total training timesteps). According to the results, SG achieves the lowest FID and highest HPS score with the shift scale set to 30.
@@
One can also dynamically adjust $\delta(t)$ during the reverse diffusion process. At higher noise levels, a large shift of $\delta(t)$ can be adopted, as noisy artifacts can be rapidly removed at the early stage of diffusion sampling. Thus, we introduce $\delta(t)$ as a linear function of time ($\delta(t)=t/\sigma$), allowing stronger guidance at higher noise levels while reducing it at lower levels. More experiment results are shown in the Appendix A.4. We find that dynamic shift scaling improves quality, avoiding blurred images or noise issues caused by excessive guidance as the noise level decreases to zero.
6 Conclusions and Limitations
This paper suggests that the reason for artifacts in image generation is incompetent denoising at each step, and proposes Self-Guidance (SG), an inference-guidance strategy to modulate the output distribution at each denoising step via incorporating the difference between the output probability distributions computed at both current time $t$ and a noisier one $t+\delta(t)$. SG has multiple advantages, such as architecture independence, requiring no specific training, and compatibility with other well-established guidance methods. Intensive experiment results demonstrate that SG can consistently improve generation performances on multiple popular text-to-image and text-to-video models, particularly effective when addressing long-standing problems like generating high-quality human hands and bodies.
While SG requires two forward passes per sampling step, it leads to increased computational cost. To address this issue, we propose an efficient method, SG-prev, which approximates SG by directly reusing the output from the immediately previous timestep. As a result, SG-prev introduces no additional computational cost.
Supplementary Material
Appendix A Additional results
Figure 7: SG results of different guidance scales and shift scales on Stable Diffusion 3-medium.
A.1 More results of SG combining with CFG and PAG
Figure 8 shows the comparison between CFG, CFG+PAG, and CFG + PAG+SG. The results show that adding SG to CFG and PAG significantly enhances the model’s ability to generate fine details (e.g., hands are rendered more realistically), and improves prompt alignment, ensuring generated images better match the intended descriptions. SG also effectively corrects errors seen in previous samples, such as extra hands or incorrect body configurations in humans and animals.
Figure 9: Ablation study of the dynamic shift scale on EDM2-XXL[24].
A.2 More results of Text-to-Video Generation
Figure 10 shows more results of our SG on Test-to-Video models, similar to Figrue 10. For complete video examples, see the videos folder of the Appendix.
Figure 10: Additional results for using SG on Text-to-Video models, similar to Figure 10
A.3 More results of ablations of guidance scale and shift scale
Figure 7 shows the results of different guidance scales and shift scales on Stable Diffusion 3-medium. We can find that with the increase of the guidance scale and shift scale, our SG can significantly improve the guidance performance.
A.4 More results of ablations of dynamic shift scale
Figure 9 shows the result of the dynamic shift scale on the EDM2-XXL model. $\sigma$ is a fixed value. We let the shift scale $\delta(t)$ be $-\sigma$, $\sigma$, and $t/\sigma$. The results show that using fixed $\delta(t)$ caused either blurred images or excessive noise, while dynamic $\delta(t)$ is better.
Appendix B Details of the flow models
B.1 Flow Matching (FM)
Flow Matching (FM) is based on continuous normalizing flows, where the generative model is defined as an ordinary differential equation (ODE):
$\frac{dz_{t}}{dt}=v(z_{t},t)$ (B.1)
Here, $t\in[0,1]$, and $v(z_{t},t)$ is referred to as the vector field. This ODE defines a probability path $p_{t}$, transitioning from a noise distribution $p_{1}$ to the data distribution $p_{0}$. Once $v(z_{t},t)$ is known, we can use an ODE solver to generate data samples.
The vector field $v_{\theta}(z_{t},t)$ is parameterized by a neural network, and the FM objective is:
$L_{FM}=\mathbb{E}_{t,p_{t}(z)}\|v_{\theta}(z,t)-u_{t}(z)\|^{2}_{2}$ (B.2)
Here, $u_{t}(z)$ is the target vector field, defining the probability path $p_{t}(z)$ from $p_{1}$ to $p_{0}$. Without priors, $u_{t}(z)$ is unknown. FM introduces a predefined $u_{t}(z)$ using a conditional probability path $p_{t}(z|x_{0})$, where $x_{0}$ is the real data. This is modeled as:
$p_{t}(z|x_{0})=\mathcal{N}(z|a_{t}x_{0},b_{t}^{2}I)$ (B.3)
Here, $a_{t}$ and $b_{t}$ are time-dependent functions. If $a_{0}=1,b_{0}=0$, then $p_{0}(z|x_{0})=q(x_{0})$, and if $a_{1}=0,b_{1}=1$, then $p_{1}(z|x_{0})=p_{1}$. This aligns with the diffusion process $z_{t}=a_{t}x_{0}+b_{t}\epsilon$, where $\epsilon\sim\mathcal{N}(0,I)$.
B.2 Conditional Flow Matching (CFM)
The Conditional Flow Matching (CFM) objective is:
$L_{CFM}=\mathbb{E}_{t,q(x_{0}),p_{t}(z|x_{0})}\|v_{\theta}(z,t)-u_{t}(z|x_{0})\|^{2}_{2}$ (B.4)
Here, $u_{t}(z|x_{0})$ defines the conditional probability path $p_{t}(z|x_{0})$. It can be derived as:
$u_{t}(z|x_{0})=\frac{a_{t}^{\prime}}{a_{t}}z_{t}-b_{t}^{2}\lambda_{t}^{\prime}\epsilon$ (B.5)
where $\lambda_{t}=\log\frac{a_{t}^{2}}{b_{t}^{2}}$ is the Signal-to-Noise Ratio(SNR), then we ca get $\lambda_{t}^{\prime}=2\left(\frac{a_{t}^{\prime}}{a_{t}}-\frac{b_{t}^{\prime}}{b_{t}}\right)$. The final CFM objective becomes:
$L_{CFM}=\mathbb{E}_{t,q(x_{0}),p_{t}(z|x_{0}),\epsilon\sim\mathcal{N}(0,I)}\left\|v_{\theta}(z,t)-\left(\frac{a_{t}^{\prime}}{a_{t}}z_{t}-b_{t}^{2}\lambda_{t}^{\prime}\epsilon\right)\right\|^{2}_{2}$ (B.6)
By defining $v_{\theta}(z,t)=\frac{a_{t}^{\prime}}{a_{t}}z_{t}-b_{t}^{2}\lambda_{t}^{\prime}\epsilon_{\theta}(z,t)$, this becomes equivalent to noise prediction like DDPM, with an additional weight term.
B.3 Rectified Flow (RF)
RF simplifies the forward process to:
$z_{t}=(1-t)x_{0}+t\epsilon$
This linear interpolation allows larger step sizes during sampling. Its loss is straightforward:
$L_{RF}=\mathbb{E}_{t,q(x_{0}),p_{t}(z|x_{0}),\epsilon\sim\mathcal{N}(0,I)}\|v_{\theta}(z,t)-(\epsilon-x_{0})\|^{2}_{2}$
For RF, $w_{t}=\frac{t}{1-t}$, showing its connection to FM and other diffusion models.
Appendix C Implementation details
All experiments are conducted on NVIDIA GeForce RTX V100 GPU and NVIDIA RTX A100 GPU for image and video sampling. We obtain all the necessary weights from publicly available repositories.
C.1 Metrics for Text-to-Image Generation
To evaluate the performance of Text-to-Image models, we employed several metrics:
  • Fréchet Inception Distance (FID) [10] quantifies the similarity between generated images and real images.
  • Human Preference Score (HPS v2.1) [64] assesses the quality of generated images based on human judgment.
  • CLIP Score [70] measures the alignment between generated images and their textual descriptions.
  • Aesthetic Score [55] evaluates the visual quality and appeal of the images.
FID, CLIP Score, and Aesthetic Score were conducted using 5000 prompts in the MS-COCO 2017 validation dataset. HPS v2.1 used the standard prompts they provided.
Stable-Diffusion 1.4 Stable-Diffusion 3-Medium Stable-Diffusion 3.5-Large Flux.1-dev
CFG guidance scale 7.5 3.5 3.5 3.5
PAG guidance scale 0.3 0.7 - -
SG guidance scale 3 3 3 3
SG constant shift scale 10 10 10 10
Table 4: The parameters of CFG, PAG, and SG used during inference for four diffusion models.
C.2 Metrics for Text-to-Video Generation
To evaluate the performance of Text-to-Video models, we utilized VBench[19], a comprehensive benchmark specifically designed for video generative. We evaluate model performance across five key dimensions: human action, scene, multiple objects, appearance style, and overall consistency.
  • Human Action: Use UMT [30] to assess if generated videos accurately depict the specific actions described in text prompts.
  • Scene: Use Tag2Text [18] to caption generated scenes and verify their consistency with the text prompt (e.g., "ocean" vs. "river").
  • Multiple Objects: Evaluate the success rate of generating all objects specified in the text prompt within each video frame.
  • Appearance Style: Use CLIP [49] to feature similarity between frames and style descriptions (e.g., "oil painting" or "cyberpunk").
  • Overall Consistency: Use ViCLIP [63] to calculate overall video-text consistency, reflecting semantic and style alignment.
C.3 Guidance Scale Setting
The guidance scales for CFG and PAG are kept the same as the original settings in corresponding papers. For each diffusion model, we provide the best guidance scale and constant shift scale of Self-Guidance. The specific parameters for these three guidance methods are in Table 4.

### v5

@@ V Experiments
TABLE II: Quantitative results of SG Performances in Specific Generative Domains compared with CFG baseline Flux.1 [29].
@@
Flux.1[29] 67.0163 0.8902 34.1976 4.5448 0.3080 29.74 22.9802 1.0971
+SG 66.4360 0.9283 34.0561 4.6634 0.3084 28.59 22.9960 1.1046
V-A Implementation Details
@@
We apply our Self-Guidance on current state-of-the-art text-to-image and text-to-video models. For text-to-image, we consider both a classic diffusion model, Stable Diffusion 1.4 [51], and current state-of-the-art flow-based models, Stable-Diffusion 3 series [8, 2] and Flux.1-dev[29]. For text-to-video, we consider CogVideoX-5B [66], an open-source video generation model trained with flow matching.
To facilitate a fair comparison, we consistently evaluate the model with the same number of steps in inference: DDIM Solver with 50 steps for SD 1.4, Euler with 28 steps for SD 3 series, and FLUX. The guidance scales for CFG and PAG are kept the same as the original settings in the corresponding papers. We will elaborate in Appendix -C3.
@@
V-B Quantitative Comparisons
V-B1 Text-to-Image Generation
@@
SG is specifically designed to address the artifact issues in CFG and PAG. As shown in Table I, incorporating SG with CFG leads to consistent improvements in HPS across all models, demonstrating that SG better aligns with human preferences. Moreover, our approach remains competitive with PAG in both CLIP and Aesthetic Score.
Then, we find that combining SG with CFG and PAG will have the best performance over using them alone. Figure 7 shows the comparison between CFG, CFG+PAG, and CFG + PAG+SG. The results show that adding SG not only enhances fine-grained details, prompt alignment, and error correction but also achieves new SoTA better performance across multiple metrics (24.65 HPS v2.1 for SD 1.4, 25.28 FID for SD 3, 0.3132 CLIP Score in SD 3.5, and 5.7845 Aesthetic Score in Flux).
In addition, SG-prev outperformed both compared baselines and the SG model, particularly in terms of FID and CLIP Score. This is achieved by applying SG-prev after the diffusion time (not step) $t$ decreases below 500, according to Figure 5. This demonstrates that the SG-prev successfully boosts the diffusion models without incurring any additional cost.
V-B2 Text-to-Video Generation
[…]
V-C Guidance Performances in Specific Domains
Through our experiments and analysis (e.g., Figure 3), we observed that SG is particularly effective when handling some long-standing problems in image generation. More specifically, SG can effectively generate hands with an exact correct number of fingers with natural shapes (Part 1), and remove artifacts like redundant arms or hands (Part 2). It also enhances Flux’s comprehension of textual input (Part 4), effectively eliminating extraneous objects. This allows the main subject of the image to stand out while the background is appropriately blurred(Part 3).
@@ V-C Guidance Performances in Specific Domains
As shown in Table II, SG achieves significant results in fine-grained generation tasks, such as hand and face generation, with improvements in FID-H (67.0163 to 66.4360) and FID-F (34.1976 to 34.0561). Additionally, the increase in PickScore and ImageReward indicates that our SG is more adept at producing higher-quality samples, thus minimizing generation errors, as illustrated in Part 4 of Figure 3.
@@
TABLE III: Quantitative results of CFG based text-to-video generation model CogVideoX with our SG and SG-prev. All scores are evaluated on the Standard Prompt Suite of VBench[19]. Bolded values highlight the best performance for each metric.
[…]
CogVideoX-5B[66] 0.8320 0.2526 0.4619 0.2331
+ SG 0.8450 0.2553 0.4703 0.2339
+SG-prev 0.8376 0.2530 0.4774 0.2333
Figure 9: Ablation study of the Guidance Scale and Shift Scale.
V-D Ablation Studies
V-D1 Guidance Scale
@@ V-D2 Shift Scale
The shift scale $\delta(t)$ represents the extent to which the noise level deviates from the current level. A simple way is to set it to a constant independent of $t$. The larger the shift scale is, the more it can contrast with the current noise level to suppress the artifacts. However, too large a shift may damage the desired patterns that should be preserved in the current sampling density, and thus a suitable shift scale should be adopted. Figure 9 shows how the FID and HPS score change as the shift scale increases from 0 to 50 (within 1,000 total training timesteps). According to the results, SG achieves the lowest FID and highest HPS score with the shift scale set to 30.
@@
One can also dynamically adjust $\delta(t)$ during the reverse diffusion process. At higher noise levels, a large shift of $\delta(t)$ can be adopted, as noisy artifacts can be rapidly removed at the early stage of diffusion sampling. Thus, we introduce $\delta(t)$ as a linear function of time ($\delta(t)=t/\sigma$), allowing stronger guidance at higher noise levels while reducing it at lower levels. More experiments are shown in Appendix -A4. We find that dynamic shift scaling improves quality, avoiding blurred images or noise issues caused by excessive guidance as the noise level decreases to zero.
V-D3 Hyperparameter Selection Strategy
Based on our experiments, setting the guidance scale ($\omega$) to 3 generally yields the best results across different models. However, slight tuning may still be necessary depending on the specific model. If the generated images appear overly blurry after setting $\omega$ = 3, increasing $\omega$ moderately (e.g., to 3.5–4) can help enhance sharpness. Conversely, if excessive noise is observed, decreasing $\omega$ (e.g., to 1–2) is recommended.
For the shift scale ($\delta(t)$), we use a default value equal to 1% of the model’s total diffusion time (e.g., $\delta(t)$ = 10 if diffusion time = 1000). The tuning strategy for $\delta(t)$ follows the same principle as for $\omega$: increase $\delta(t)$ (e.g., to 10-20) if images are too blurry, or decrease it (e.g., to 1–10) if observe excessive noise.
That said, we recommend using SG-prev as a more practical alternative, since it can be viewed as a form of dynamic shift scale. With SG-prev, only the guidance scale needs to be tuned, simplifying hyperparameter selection.
VI Conclusions
This paper suggests that the reason for artifacts in image generation is incompetent denoising at each step, and proposes Self-Guidance (SG)—an inference-time strategy that modulates the output distribution using the difference between distributions at the current step $t$ and a noisier step $t+\delta(t)$. SG is architecture-agnostic, training-free, and compatible with existing guidance methods. Extensive experiments show that SG consistently improves text-to-image and text-to-video generation, especially in challenging cases like generating realistic human hands and bodies. To reduce SG’s two forward passes per sampling step, we introduce SG-prev, which approximates SG by directly reusing the output from the immediately previous timestep. As a result, SG-prev introduces no additional computational cost.